# Security and Real-time FPGA integration for Learned Image Compression

### v1

@@ Abstract
Learnable Image Compression (LIC) has proven capable of outperforming standardized video codecs in compression efficiency. However, achieving both real-time and secure LIC operations on hardware presents significant conceptual and methodological challenges. The present work addresses these challenges by providing an integrated workflow and platform for training, securing, and deploying LIC models on hardware. To this end, a hardware-friendly LIC model is obtained by iteratively pruning and quantizing the model within a standard end-to-end learning framework. Notably, we introduce a novel Quantization-Aware Watermarking (QAW) technique, where the model is watermarked during quantization using a join loss function, ensuring robust security without compromising model performance. The watermarked weights are then public-key encrypted, guaranteeing both content protection and user traceability. Experimental results across different FPGA platforms evaluate real-time performance, latency, energy consumption, and compression efficiency, highlighting that the watermarking and encryption processes maintain negligible impact on compression efficiency (average of -0.4 PSNR) and energy consumption (average of +2%), while still meeting real-time constraints and preserving security properties.
@@ I Introduction
In Learned Image Compression (LIC) [1, 2, 3, 4], an autoencoder is trained to encode and decode an image while generating a compressed bitstream in the process [5, 6]. At the source, the image is projected into a low-dimensional latent space by the encoder model. This representation is then quantized and entropy-coded into a binary bitstream. At the receiver, the bitstream is entropy-decoded, and the decoder model projects the representation back to the pixel domain, recovering an approximation of the original image. Over the last years, the LIC efficiency has seen consistent advancements, nowadays challenging standardized coding technologies and consequently, the interest in its realtime hardware-based LIC implementations emerged.
Research on realtime LIC for hardware has received lot of attention, tackling challenges such as computational/storage/power efficiency [7, 8], cross-platform support [9], and compatibility with edge computing devices [10, 11]. Achieving real-time operations on hardware usually relies a combination of strategies. The model is first pruned to minimize the number of filters or neurons [12, 13]. Then, the surviving parameters are quantized [14] to cope with the fixed-point computation capabilities of embedded hardware. For example, [15] and [16] propose a FPGA implementations of LIC models with 8-bit parameter quantization. The state of the art architecture [17] achieves 37 fps encoding and decoding at 720p resolution on a KU115 FPGA.
@@
Neural network watermarking consists in hiding an imperceptible fingerprint within a model parameters at training time, so that the model integrity can be verified later on [18]. Watermarking a LIC models requires that the watermark does not alter the video quality or encoding efficiency. Watermarking a model deployed on hardware [19] is even more challenging due to the constraint of hiding the watermark within the few parameters surviving the pruning step. However, attacks aside, the quantization step alone tends to destroy the watermark that cannot be easily retrieved anymore. Atop of that, retrieving and verifying the watermark must not compromise realtime operations. Therefore, securing a hardware based LIC model while preserving its properties is a significant challenge that has received little attention to date and is the object of this work.
@@ II-B Neural Network Watermarking
While most of the watermarking literature focuses on classification tasks, works demonstrate that white-box method can be extended beyond this realm [28, 29]. Various computer vision models have been successfully watermarked against a wide range of attacks [30, 25]. Only [19] apply watermarking at a hardware-level but still needs to bridge the gap between algorithmic design and efficient hardware implementation and to reach the trade-off among security, latency, and impact on video quality. Yet, no papers tackle the challenge of applying watermarking to a real-quantized LIC model. The present study aims to address the challenges of watermarking learned image compression (LIC) neural networks: Can a watermark be embedded in a quantized model without impacting inference time or degrading image quality?
@@ IV-A Training the Model
We preliminary train from scratch the LIC model [5] to validate our experimental setup and to collect baseline numbers for compression efficiency and latency. As per standard practice, we train 5 different instances of the model times for the $\lambda$ values (0.001, 0.005, 0.01, 0.05, 0.1) to span a large enough range of RD trade-offs. In detail, we consider a hardware-friendly implementation and training procedure of the model in Fig. 1 inspired by related literature to [15] where the GDN activation is replaced with a simple ReLU to cut latency towards realtime operations. While [17] achieves higher system-level performance on the same LIC model, that is at the cost of a complex hardware design space exploration that is out of the scope of this work. The model is trained using the same randomized train/validate/test dataset split of the combined CLIC, DIV2K, and a subset of the Open Images datasets. Fig. 6 shows the encoding efficiency of our baseline, hardware-friendly, implementation of [5] versus the original model. The efficiency drop due to the use of the simpler ReLU in place of the GDN and is in line with what observed in similar literature. We speculate that a hardware-optimized implementation of the GDN may preserve encoding efficiency while minimizing the impact on latency, however this is outside the goals of the present work.
@@
Table I presents the encode-and-decode latency for different image sizes in a manner consistent with related work [15], measured on a Tesla K80 GPU. Following the approach in [15], we generate HD (1280x720), FHD (1920x1080), and UHD (3840x2160) images by repeatedly tiling images from the aforementioned datasets until the target resolution is achieved. To account for the memory constraints typical of embedded platforms, the image is processed as partially overlapping 256x256 patches from the tiled images, as in related literature. Each patch is independently encoded and decoded, and the decoded patches are finally reassembled into the whole picture. The table shows that the reference LIC model cannot achieve realtime video coding despite our optimizations, emphasizing the need for hardware-tuned LIC models especially for high-resolution contents.
TABLE I: Model-specific performance on a single Tesla K80 GPU. Up/down arrows always indicate higher/lower is better.
@@
Latency (s) $\downarrow$ 2.88 6.7 26.4
@@ IV-B Pruning the Model
IV-C Quantizing the Model
TABLE III: Effect of Quantization Aware Training (QAT) [24] compared to the pruned float model
HD FHD UHD
$\Delta$ Model size ↓ -42.86% -48.85% -48.85%
$\Delta$ Inference latency ↓ -2.56% -3.25% -4.12%
$\Delta$ Avg PSNR (dB) ↑ +0.130 +0.092 +0.11
$\Delta$ Avg MS-SSIM (dB) ↑ +0.110 +0.071 +0.086
As a second preliminary step, we quantize the pruned model weights using the same Quantization-Aware Training (QAT) process from [24], without adding any watermark. This experiment is meant to isolate the effects of quantization alone, establishing a baseline for assessing the impact of watermarking in the proposed QAW strategy later on. We run QAT for a 1000 iterations using a separate training subset sampled from the original datasets. Table III shows the changes in model size, accuracy, and inference latency across HD, FHD, and UHD resolutions. Quantization further reduces the model size down to nearly 50%, making the model more memory-efficient for deployment on resource-constrained devices. This compression is accompanied by a modest gain in inference latency (-2.56% to -4.12%), improving real-time performance without substantial sacrifices in speed. Interestingly, unlike pruning, quantization results in slight improvements in both PSNR (up to +0.130 dB) and MS-SSIM (up to +0.110). We hypothesize that quantization noise acts akin to a regularizer, preventing overfitting on the training data and improving generalization over data unseen at training time.
IV-D Watermaking the Model
In this section, we finally experiment with hiding a watermark within the pruned model weights during the quantization step with our novel QAW strategy as described in Sec. III-B. Prior to training, we extract the minimum and maximum values from various model layers. These values allow the quantizer to derive the optimal scale and zero-point for the floating-point model, based on the calibration subset we provide. The calibration set for all quantization experiments consists of 1000 randomly selected images from CLIC [37], DIV2K [38], and Open Images [39], which are used exclusively for the forward pass. For the actual quantization training, we run 1000 iterations using a randomly selected subset of 5000 images from CLIC, DIV2K, and Open Images. Symmetric quantization is employed with the zero-point set to zero and saturation enabled, minimizing quantization error. We set $\beta=0.1$ and $\lambda$ to the same values used during the floating-point model training, allowing us to explore different rate-distortion (RD) trade-offs ($0.001,0.005,0.01,0.05,0.1$). Next, we evaluate the impact of the introduced watermarking on the model through the joint QAW training scheme. We measure the robustness of the watermark in C-BER terms [28], which denotes the differences between the embedded and the extracted watermarks expressed in terms of the bit error rate (BER), as follows:
@@ IV-D Watermaking the Model
The lower such metric, the easier the embedded watermarking can be recovered from the model. In detail, the watermark is detected by projecting the watermarked layer onto the secret key, this is done on the FPGA PS using the runtime API. The results are reported in Table IV and show that the watermark is successfully retrieved in almost all cases for a negligible drop in image quality with respect to the float model.
@@
TABLE IV: Effect of the proposed Quantization Aware Watermarking (QAW) at $\lambda=0.01$ and $\beta=0.01$ compared to the pruned float model
HD FHD UHD
$\Delta$ Avg PSNR (dB) ↑ -0.35 -0.19 -0.25
$\Delta$ Avg MS-SSIM (dB) ↑ -0.17 -0.091 -0.12
C-BER ↑ 95% 98% 96%
@@
To evaluate the efficiency of QAW, we consider a simple Post Quantization Watermarking (PQW) technique where the watermark is embedded in the quantized model [40]. In detail, a parameter tensor $T$ from the host network is flattened, trimmed to length $N^{2}$, and reshaped into an $N\times N$ matrix $T_{\text{sq}}$. The watermark, generated using a key $k$, is embedded by shuffling elements, applying the discrete cosine transform (DCT), and modifying the mid-frequency DCT coefficients. After embedding, $T_{\text{sq}}$ is resized to match $T$, and the host network’s parameters are updated. During extraction, the process is reversed: the mid-frequency domain of $T_{\text{sq}}$ is extracted, and pseudo-random numbers are generated using $k$. The correlation between the extracted mid-frequency domain and the generated pseudo-random numbers is calculated. A strong correlation indicates the presence of the watermark. We target the first convolutional layer and set and use the same setup variables from [40] The model is then evaluated again on the test subset and the changes in performance are reported in Table V. While this scheme demonstrates slightly improved robustness, its adverse effect on reconstructed image/video quality is higher than that of QAW, as expected.
@@
TABLE V: Performance changes due to the reference Post-Quantization Watermarking (PQW) [40] compared to the pruned float model
@@
Our experiments demonstrated a trade-off between watermark robustness and impact on image quality. Since both methods ensure watermark presence with comparable robustness, we will only consider the QAW scheme, as it results in an average of 0.3 PSNR (dB) improvement over PQW, proving to be the better approach.
@@
TABLE VI: Power consumption comparison with related works and GPU [J/frame]
@@
TABLE VII: Resource utilization for HD LIC using B4096 DPU cores and comparison with state of the art
@@
TABLE VIII: Impact of DRM on FPS and Energy (J/frame) for KV260. DRM refers to the combination of watermark-extraction and decryption applied to the baseline model.
Resolution Baseline Baseline + DRM $\Delta$ Energy (%) $\Delta$ FPS (%)
FPS Power [J/frame] FPS Power [J/frame]
HD (1280x720) 26.8 FPS 3.00 J/frame 26.4 FPS 3.12 J/frame +4.0% -1.5%
FHD (1920x1080) 9.9 FPS 8.40 J/frame 9.2 FPS 8.59 J/frame +2.3% -7.1%
UHD (3840x2160) 4.8 FPS 31.50 J/frame 4.2 FPS 31.71 J/frame +0.7% -12.5%
TABLE IX: Impact of DRM on FPS and Energy (J/frame) for ZCU102. DRM refers to the combination of watermark-extraction and decryption applied to the baseline model.
Resolution Baseline Baseline + DRM $\Delta$ Energy (%) $\Delta$ FPS (%)
FPS Power [J/frame] FPS Power [J/frame]
HD (1280x720) 62.5 FPS 2.80 J/frame 61.2 FPS 2.88 J/frame +2.9% -2.1%
FHD (1920x1080) 25.5 FPS 5.30 J/frame 24.2 FPS 5.44 J/frame +2.6% -5.1%
UHD (3840x2160) 15.8 FPS 19.80 J/frame 14.3 FPS 20.23 J/frame +2.2% -9.5%
IV-E On-board experiments
TABLE X: FPS performance for the proposed hardware model using KV260 and ZCU102 FPGAs

### v2

@@ Abstract
Learned Image Compression (LIC) has demonstrated superior compression efficiency compared to standardized video codecs. However, achieving both real-time performance and robust security in hardware-based LIC deployments remains a significant challenge. This work addresses these challenges by presenting an integrated workflow and deployment platform for training, securing, and implementing LIC models on hardware. To achieve a hardware-efficient LIC model, we employ an iterative pruning and quantization process within a standard end-to-end learning framework. Additionally, we introduce Quantization-Aware Watermarking (QAW)—a novel technique that embeds a watermark directly during quantization via a joint loss function. This ensures model integrity and security without degrading performance. To further enhance protection, the watermarked weights undergo public-key encryption, safeguarding both content and user traceability. We evaluate real-time performance, latency, energy consumption, and compression efficiency across multiple FPGA platforms. Results indicate that the watermarking and encryption steps introduce only a minor overhead—PSNR decreases by 0.2 dB on average, energy consumption increases by 2%, and FPS drops by 6%—while maintaining real-time constraints and security properties. Furthermore, our approach outperforms existing hardware-based LIC implementations in both FPS and energy efficiency, delivering optimized LIC codecs for HD, FHD, and UHD resolutions.
@@ I Introduction
In Learned Image Compression (LIC) [1, 2, 3, 4], an autoencoder is trained to encode and decode an image while generating a compressed bitstream in the process [5, 6]. At the source, the image is projected into a low-dimensional latent space by the encoder model. This representation is then quantized and entropy-coded into a binary bitstream. At the receiver, the bitstream is entropy-decoded, and the decoder model projects the representation back to the pixel domain, recovering an approximation of the original image. Over the last years, the LIC efficiency has seen consistent advancements, nowadays challenging standardized coding technologies [7, 8] and consequently, the interest in its realtime hardware-based LIC implementations emerged.
Research on realtime LIC for hardware has received lot of attention, tackling challenges such as computational/storage/power efficiency [9, 10], cross-platform support [11, 12], and compatibility with edge computing devices [13, 14, 15]. Achieving real-time operations on hardware usually relies a combination of strategies. The model is first pruned to minimize the number of filters or neurons [16, 17]. Then, the surviving parameters are quantized [18] to cope with the fixed-point computation capabilities of embedded hardware. For example, [19] and [20] propose a FPGA implementations of LIC models with 8-bit parameter quantization. The state of the art architecture [21] achieves 37 fps encoding and decoding at 720p resolution on a KU115 FPGA.
@@
Neural network watermarking [22] consists of hiding an imperceptible fingerprint within a model parameters at training time, so that the model integrity can be verified later on [23]. Watermarking a LIC models requires that the watermark does not alter the video quality or encoding efficiency and survives potential attacks [24]. Watermarking a model deployed on hardware [25] is even more challenging due to the constraint of hiding the watermark within the few parameters surviving the pruning step. However, attacks aside, the quantization step alone tends to destroy the watermark that cannot be easily retrieved anymore. Atop of that, retrieving and verifying the watermark must not compromise realtime operations. Therefore, securing a hardware based LIC model while preserving its properties is a significant challenge that has received little attention to date and is the object of this work.
@@ II-B Neural Network Watermarking
While most of the watermarking literature focuses on classification tasks, recent works demonstrate that white-box methods can extend beyond this domain [34, 35]. Various computer vision models have been successfully watermarked against a wide range of attacks [36, 31]. However, only [25] applies watermarking at the hardware level, yet it still faces challenges in bridging the gap between algorithmic design and efficient hardware implementation while balancing security, latency, and video quality. Notably, no prior work addresses the challenge of watermarking a real quantized Learned Image Compression (LIC) model. This study aims to explore the difficulties of LIC neural network watermarking, posing the following key questions: Can a watermark be inserted into a quantized model? Can this insertion be achieved without significant impact on latency and energy or degrading image quality?
@@ IV-A Training the Model
We preliminarily train the LIC model [5] from scratch to validate our experimental setup and establish baseline metrics for compression efficiency and latency. Following standard practice, we train five different instances of the model for various values of $\lambda$ (0.001, 0.005, 0.01, 0.05, 0.1) to cover a sufficiently wide range of rate-distortion (RD) trade-offs. Specifically, we consider a hardware-friendly implementation and training procedure of the model, as shown in Fig. 1, which is inspired by related work such as [19]. In this approach, the Generalized Divisive Normalization (GDN) layer activation is replaced with a simpler ReLU activation to reduce latency and streamline the implementation for real-time operations. While [21] achieves higher system-level performance on the same LIC model, that is at the cost of a complex hardware design space exploration that is out of the scope of this work. The model is trained using the same randomized train/validate/test dataset split of the combined CLIC, DIV2K, and a subset of the Open Images datasets. Fig. 6 illustrates the encoding efficiency of our baseline, hardware-friendly implementation of [5] compared to the original model. The model presented in [21] is based on the computationally heavier architecture from [6], which incorporates an additional hyperprior network instead of a simple factorized entropy encoder. This modification leads to a noticeable improvement in rate-distortion (RD) performance, albeit at the cost of system-level performance. The observed efficiency drop due to the use of a simpler ReLU activation function in place of the GDN is consistent with findings in similar literature. We hypothesize that a hardware-optimized implementation of the GDN could potentially preserve encoding efficiency while minimizing latency impact; however, exploring this is beyond the scope of the present work.
@@
Table I presents the encode-and-decode latency for different image sizes in a manner consistent with related work [19], measured on a A100-PCIE-40GB GPU. Following the approach in [19], we generate HD (1280x720), FHD (1920x1080), and UHD (3840x2160) images by repeatedly tiling images from the aforementioned datasets until the target resolution is achieved. To account for the memory constraints typical of embedded platforms, the image is processed as partially overlapping 256x256 patches from the tiled images, as in related literature. Each patch is independently encoded and decoded, and the decoded patches are finally reassembled into the whole picture. The table shows that the reference LIC model cannot achieve realtime video coding despite our optimizations, emphasizing the need for hardware-tuned LIC models especially for high-resolution contents.
TABLE I: Model-specific performance on a single A100-PCIE-40GB GPU. Up/down arrows indicate higher/lower is better.
@@
Latency (s) $\downarrow$ 0.35 0.79 3.16
Avg PSNR (dB) $\uparrow$ 33.95 33.83 32.91
Avg MS-SSIM $\uparrow$ 0.982 0.978 0.967
@@ IV-B Pruning the Model
IV-C Jointly Watermarking and Quantization the Model
In this section, we experiment with hiding a watermark within the pruned model weights during the quantization step with our novel QAW strategy as described in Sec. III-B. Prior to training, we extract the minimum and maximum values from various model layers. These values allow the quantizer to derive the optimal scale and zero-point for the floating-point model, based on the calibration subset we provide. The calibration set for all quantization experiments consists of 1000 randomly selected images from CLIC [43], DIV2K [44], and Open Images [45], which are used exclusively for the forward pass.
For the actual quantization training, we run 1000 iterations using a randomly selected subset of 5000 images from CLIC, DIV2K, and Open Images. Symmetric quantization is employed with the zero-point set to zero and saturation enabled, minimizing quantization error. We set $\lambda$ to the same values used during the floating-point model training, allowing us to explore different rate-distortion (RD) trade-offs ($0.001,0.005,0.01,0.05,0.1$). As for the watermark embedding loss function variable $\beta$, we explore a range of values based on the chosen $\lambda$ and target layer in order to find the optimal $\lambda$ $\beta$ pair for each model. To achieve this, we adaptively modify $\beta$ during QAW to ensure that the watermark is fully embedded by the end of the joint-training, in the event of failure to embed, $\beta$ is raised, and if embedding is achieved quickly $\beta$ is lowered.
TABLE III: Performance metrics for pre-watermarked model.
Lambda PSNR MSSIM BPP
0.1 38.73 0.992 1.567
0.05 37.72 0.9899 1.205
0.01 33.57 0.975 0.440
0.005 31.81 0.9627 0.261
0.001 28.12 0.921 0.092
Once QAW is done, we evaluate the impact of the introduced watermarking on the model by measuring the BD metrics using the float-pruned model as a reference, following the joint QAW training scheme. Table III presents the PSNR, MSSIM, and BPP values that serve as the baseline for our analysis. To assess the effect of watermarking, we experiment with embedding the watermark across all four layers of the decoder.
We evaluate the robustness of the watermark in terms of C-BER [34], which quantifies the differences between the embedded and extracted watermarks, expressed in terms of the bit error rate (BER), as follows:
@@ IV-C Jointly Watermarking and Quantization the Model
In this case, a higher C-BER value indicates better robustness of the watermark, as it reflects a smaller error between the embedded and extracted watermarks, signifying a more reliable watermark. The watermark detection is then performed by projecting the watermarked layer onto the secret key, a process executed on the FPGA processing system using the runtime API.
The experimental results are presented in Tables VII, VI, V, and IV. These tables display the PSNR, MSSIM, and BPP performance of the five trained models with the watermark embedded in different layers.It is observed that embedding the watermark in the final layer is particularly challenging, a phenomenon that has been well-documented in the neural network watermarking literature [34]. This difficulty stems from the final layer’s heightened sensitivity to perturbations in the network parameters, particularly the weights and activations. Unlike earlier layers, the final layer is responsible for generating the output of the model, where even small changes to the parameters can significantly affect the model’s predictions. The final layer typically operates in a more linear regime compared to earlier layers, where activations are often saturated or subject to nonlinearities. Therefore, modifications introduced by watermark embedding in the final layer tend to have a disproportionately large effect on the model’s performance, as they directly influence the output logits.
TABLE IV: Performance metrics for Layer 4 with watermarking.
Lambda / Beta PSNR MSSIM BPP
0.1 / 0.1 38.31 0.988 1.681
0.05 / 0.1 37.31 0.981 1.275
0.01 / 0.05 33.22 0.970 0.470
0.005 / 0.02 31.51 0.960 0.290
0.001 / 0.01 27.86 0.910 0.096
TABLE V: Performance metrics for Layer 3 with watermarking.
Lambda / Beta PSNR MSSIM BPP
0.1 / 0.1 38.67 0.9924 1.640
0.05 / 0.1 37.51 0.989 1.244
0.01 / 0.05 33.54 0.975 0.480
0.005 / 0.02 31.79 0.964 0.280
0.001 / 0.01 28.09 0.920 0.0944
TABLE VI: Performance metrics for Layer 2 with watermarking.
Lambda / Beta PSNR MSSIM BPP
0.1 / 0.1 38.69 0.993 1.645
0.05 / 0.1 37.57 0.991 1.211
0.01 / 0.05 33.61 0.976 0.462
0.005 / 0.02 31.83 0.964 0.284
0.001 / 0.01 28.12 0.921 0.0951
TABLE VII: Performance metrics for Layer 1 with watermarking.
Lambda / Beta PSNR MSSIM BPP
0.1 / 0.1 38.54 0.992 1.679
0.05 / 0.1 37.58 0.990 1.206
0.01 / 0.05 33.52 0.976 0.465
0.005 / 0.02 31.74 0.964 0.298
0.001 / 0.01 28.13 0.951 0.094
Our experiments corroborate this observation, as watermarking the final layer resulted in the highest performance degradation across all layers. The embedding difficulty in the final layer translates into higher $\beta$ values in the QAW loss function, reflecting a stronger penalty for watermarking during training. Consequently, this increase in $\beta$ leads to a conflict with the rate-distortion (RD) loss, ultimately resulting in a notable reduction in RD performance due to the competing objectives of watermark preservation and model accuracy. Furthermore, it is evident that the $\beta$ value decreases with lower $\lambda$ values. This is intuitive, as the QAW scheme jointly minimizes both loss terms, driven by the tradeoff variables $\lambda$ and $\beta$. When $\lambda$ is reduced, watermarking becomes easier, leading to an adaptive decrease in $\beta$. Finally, the QAW results are summarized in Fig.7 and Table VIII, which further highlight the insights discussed above. Overall, Layer 2 proves to be the optimal location for watermarking, yielding the lowest RD performance loss (-0.2 dB and +5.11% BPP) and demonstrates that the watermark is successfully recovered in nearly all cases, with only a negligible degradation in image quality compared to the float model. In addition, quantization further reduces the model size down to nearly 43%, making the model more memory-efficient for deployment on resource-constrained devices. This compression is accompanied by a modest gain in inference latency (-2.56%), improving real-time performance without substantial sacrifices in speed.
Figure 7: RD performance of proposed QAW with pruned float model reference [5] at five $\lambda$ $\beta$ pairs
@@
TABLE VIII: Effect of the proposed Quantization Aware Watermarking (QAW) compared to the pruned float model
Layer 1 Layer 2 Layer 3 Layer 4
$\Delta$ BD-PSNR (dB) ↑ -0.34 -0.20 -0.29 -0.62
$\Delta$ BD-Rate (%)↓ +8.07 +5.11 +8.10 +17.7
C-BER ↑ 97% 98% 98% 94%
$\Delta$ Model size ↓ -42.86%
$\Delta$ Inference latency ↓ -2.56%
@@
To evaluate the efficiency of QAW, we consider a simple Post Quantization Watermarking (PQW) technique where the watermark is embedded in the quantized model [46]. In detail, a parameter tensor $T$ from the host network is flattened, trimmed to length $N^{2}$, and reshaped into an $N\times N$ matrix $T_{\text{sq}}$. The watermark, generated using a key $k$, is embedded by shuffling elements, applying the discrete cosine transform (DCT), and modifying the mid-frequency DCT coefficients. After embedding, $T_{\text{sq}}$ is resized to match $T$, and the host network’s parameters are updated. During extraction, the process is reversed: the mid-frequency domain of $T_{\text{sq}}$ is extracted, and pseudo-random numbers are generated using $k$. The correlation between the extracted mid-frequency domain and the generated pseudo-random numbers is calculated. A strong correlation indicates the presence of the watermark. We target the first convolutional layer and set and use the same setup variables from [46] The model is then evaluated again on the test subset and the changes in performance are reported in Table IX. While this scheme demonstrates slightly improved robustness, its adverse effect on reconstructed image/video quality is higher than that of QAW, as expected.
@@
TABLE IX: Performance changes due to the reference Post-Quantization Watermarking (PQW) [46] compared to the pruned float model
@@
Our experiments demonstrated a trade-off between watermark robustness and impact on image quality. Since both methods ensure watermark presence with comparable robustness, we will only consider the QAW scheme, as it results in an average of 0.5 PSNR (dB) improvement over PQW, proving to be the better approach.
@@
TABLE X: Example model ($\lambda$=0.01) showing changes in PSNR at different steps of the workflow evaluated on the CLIC2020 dataset
HD FHD UHD
Float-Model 34.95 34.83 33.91
Pruning 33.84 33.15 33.13
QAW 33.61 32.89 32.71
TABLE XI: Power consumption comparison with related works and GPU [J/frame]
@@
Jetson Nano (5W) 17.43 29.91 68.93
TABLE XII: Resource utilization for HD LIC using B4096 DPU cores and comparison with state of the art
@@
IV-D On-board experiments
TABLE XIII: FPS performance for the proposed hardware model using KV260 and ZCU102 FPGAs